OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 3) and (2, 7)?

OpenStudy (anonymous):

slope=m=(5-3)/(0+2)=2/2=1 y=x+b 3=-2+b, b=5 y=x+5